Status Report

NASA Space Station On-Orbit Status 14 May 2005

By SpaceRef Editor
May 15, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 14 May 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

As generally on Saturdays, Sergei Krikalev and John Phillips were off duty today, except for some housekeeping and voluntary work.

First, the crew performed the regular weekly 3-hr task of thorough station cleaning, wearing protective garment. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

Then, the CDR did the daily routine maintenance of the Service Module (SM)’s SOZh environment control & life support system, including its ASU toilet system, plus today also the weekly inspection of the BRPK air/liquid condensate separator apparatus.

The FE/SO took the periodic (weekly) reading of the cabin air’s current CO2 partial pressure in the SM and Lab, using the U.S. CDMK (CO2 monitor kit), for calldown (along with the battery status) for use in trending analyses.

At ~10:00am EDT, the crew held their weekly teleconference with ISS Program Management at JSC/Houston via S-band/audio.

At ~10:20am, Sergei and John engaged in the regular weekly planning conference (WPC) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-H and TsUP/Moscow timeline planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.

And at 11:00am, Phillips had his weekly PFC (private family conference) via S-band/audio and Ku-band/NetMeeting video.

During the monthly TVIS maintenance/inspection yesterday (5/13), the crew found the treadmill damaged: the starboard forward gyroscope wire rope is completely severed and a wire rope on the portside forward corner bracket is frayed. The crew was asked to photodocument the two findings and stand down on TVIS exercise until adequate repairs have been made. [Last night, engineers determined that the crew should go ahead with replacing the gyro wire rope on Monday (5/16), except that the only available suitable torque wrench on board has expired its certification. Thereupon, a stress test was performed at MCC-H on the gyro wire rope clamp bolts, and today the use of the torque wrench was approved for Monday’s R&R.]

The crew conducted their regular daily 2.5-hr. physical exercise program on the CEVIS cycle ergometer, RED resistive exerciser and VELO bike with bungee cord force loader (NS-1). As stated, TVIS is currently off limits.

The FE then transferred the exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of his workout on RED and CEVIS, followed by their erasure on the HRM storage medium.

Sergei Krikalev had three new voluntary work items on his “job jar” task list for today, headed off by his first run of the regular monthly sessions of the VC8 “Eneide” program’s ETD experiment (Investigation of the Coordination of Eye and Head Movements). [After a calibration with the calibrating unit, the experiment investigates horizontal eye and head movement coordination, measured Listing’s plane, and determined the orientation of the vestibulo-ocular coordinate system, using five target marks on the horizontal plane. Each step required another prior calibration run, using visual target cues or the calibration unit.]

Also off his task list, Sergei conducted a new session with the “Uragan” (hurricane) earth-imaging program, focusing the Nikon D1X No. 3 digital camera with 800-mm lens on targets called out on an uplinked list. [Targets for today were farm lands in Brazil, the western region of Lake Sevan, the Terek River valley towards the coast, Mangyshlak, and the Panama Canal.]

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

A third new item in the Russian “job jar” was the long-term recurring task of imaging the externally mounted PKZ-1V Kromka 1-3 contamination experiment tablet. [The Kromka tablet, deployed on handrail 2614 of the DC-1 “Pirs” docking compartment, collects thruster plume effluents. The pictures are taken with the Kodak 760 digital still camera (DSC) from the EVA hatch 1 (VP1) “illyuminator” (window) in the DC-1.]

Planning is underway on the ground for the crew to clean out PMA-2 (Pressurized Mating Adapter #2) in preparation for the LF-1 docking in July. About 4 hrs. of crewtime are estimated for the cleanout and transfer activities. [New location for most of the equipment will be PMA-1 (crew transfer bags, containers), and PMA-3 (for the bagged BMRRM/Bearing Motor Roll Ring Module.]

Weekly Science Update (Expedition Eleven — 3rd):

Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology (HRF GASMAP): Next week the Science Officer will conduct his first GASMAP on-orbit activity with the Routine Health Check. The HRF team looks forward to working with John.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): Nothing new.

Renal Stone (RS): The crew was thanked for the BCR (barcode reader) data of their diet intake and sample logs for the first collection session last week which “contribute greatly to the success of the experiment “. Two more sessions remain to be done. The second session will occur sometime mid-Increment after more hardware arrives on 18P or LF-1. The third session will occur late in the Increment within 30 days before the crew’s departure.

Foot/Ground Reaction Forces during Space Flight (FOOT): The SO was thanked enthusiastically for his performance of the EMG calibration dry run on real-time TV. The Foot team and PI are working to analyze the entire downlinked video and will provide Dr. Phillips with words prior to next week’s data collection.

Space Acceleration Measurement System (SAMS): SAMS remains in nominal operations.

Microgravity Acceleration Measurement System (MAMS): MAMS monitored the recent reboost. Initial MAMS measurements showed a delta-delta V of 0.68 m/s and a mean x-axis acceleration of 0.40 mg.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.

Binary Colloidal Alloy Test-3 (BCAT-3): BCAT-3 Slow Growth Sample Module will be left undisturbed in its current location by the E11 crew. In order for the samples to potentially grow crystals that can be photographed during Increment 12 operations, the Sample Module must be left undisturbed.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data. To be exchanged during LF-1.

Dust and Aerosol Measurement Feasibility Test (DAFT): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Nothing new. Will be back in July.

Serial Network Flow Monitor (SNFM): All done for Increment 10.

Fluid Merging Viscosity Measurement (FMVM): Nothing new. Will roll over to Increment 11.

Space Experiment Module (SEM): Nothing new. Experimenters and kids are working to get the next two satchels on ULF1.1.

Miscible Fluids in Microgravity (MFMG): Nothing new. Will roll over to Increment 11.

Educational Payload Operations (EPO): Nothing new. Will roll over to Increment 11.

Crew Earth Observations (CEO): As of 5/10, a total of 614 CEO images have been downlinked by Expedition 11 and reviewed on the ground. Despite the lack of light and required targets the crew was able to acquire a number of interesting features for developing their photographic techniques. Their shots include a couple of very nice passes of the rarely-photographed isles of Kerguelen in the southern Indian Ocean. CEO details of the glaciers and terrain have been judged “beautiful”. The crew also took the finest mapping set of the Falklands that investigators have seen to date from the ISS to date, and the ground meteorologist “is absolutely delighted with your long-lens high oblique views of the vertical structures of clouds”.

Today’s CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Internal waves, South China Sea-N (weather and lighting looked good at the time of this pass for sun glint enhancement of internal wave features in the South China Sea. As ISS tracked northeastward across southern Vietnam, the crew was to shoot left of track to the area between the coast and Hai-nan Island at the entrance to the Gulf of Tonkin), Pilcomayo River dynamics, N Argentina (the Pilcomayo River breaks out of the eastern Andes as a major river in southern Bolivia. However, as it flows southeastward marking the border between Argentina and Paraguay it all but vanishes into the Chaco Plain with only vegetation indicating older, abandoned channels as the river retreated. On this pass the crew used the long lens for mapping the river from where it emerges from the mountains until it became diffuse 150 miles to the SE. This pass was near the center of the target area, so the crew first looked just left of track for the lightly colored riverbed that gradually darkens with vegetation just right of track), and Plum Island Ecosystem, Maine (this is one of two Long Term Ecological Research [LTER] sites in the New England area that ISS/CEO will continue to acquire during Increment 11. [The National Science Foundation established the LTER network over twenty years ago. This network is comprised of study sites located throughout the continental United States, Alaska, Puerto Rico, Antarctica, and the Pacific Ocean, covering a wide range of ecosystems including reefs and coastal zones; hot and cold deserts; temperate, montane and grassland regions; and urban areas. The core mission of the network is to understand ecological phenomena over long temporal and spatial scales, conduct well-designed and documented experiments at and between sites, and provide information for identification and solution of ecological problems.] For this pass, the crew began with context views of southern and coastal New England. With fine weather and fairly good light, they aimed the camera left of track and began a mapping pass from the New York City area to the upper coast of Maine.

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 11 crew visit:

Expedition 11 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

  • Progress M-52 (17P) undock — 6/16;
  • Progress M-53 (18P) launch — 6/17 (dock 6/19);
  • Reboost — ~6/22 (delta-V 1.5 m/s);
  • LF-1/STS-114 launch — NET 7/13 (window opens);
  • Soyuz TMA-6 (10S) relocate (from DC-1 to FGB) — 8/16;
  • Progress M-53 (18P) undock — 8/23;
  • Progress M-54 (19P) launch — 8/24 (dock 8/26);
  • ULF1.1/STS-121 launch — NET 9/9 (window opens);
  • Soyuz TMA-7 (11S) launch — 9/27 (dock 9/29);
  • Soyuz TMA-6 (10S) return — 10/7.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at


SpaceRef staff editor.